Fermi National Laboratory


Accelerator Update for September 16, 2005 - September 19, 2005

Friday September 16
The day shift began with Operations monitoring store 4388 and the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using colliding beam, and with MiniBooNE, NuMI, and the Meson MCenter experiments taking beam.

At 8:34 AM, Operations reported that the intensity out of the 750 KeV line dropped. Over the midnight shift, operators had replaced the H- Source's command module. Operators began adjusting the source and the 750 Buncher phase to bring the intensity back up.

Operations began preparing the Recycler at 11:41 AM to stash antiprotons. The stash ended at 12:28 PM and Pbar and MCenter resumed normal operations.

Pbar experts began a reverse proton study at 1:15 PM. They completed the study at 5:24 PM.

At 6:03 PM, a NuMI beam permit tripped off due to a power supply (QF105) current problem. This happened six times in eighteen minutes. Operators changed out a control card.

At 6:51 PM, Operations reported losing the Main Injector (MI) and Recycler RF due to low LCW pressure; a LCW pump had tripped off. An operator notified the Duty Mechanic. With no RF the Recycler stash was slowly decaying away. An operator called in a RF expert.

Operators reported at 7:58 PM that the RF expert had arrived and that the RF was coming back to normal operating parameters. The Recycler had stopped losing antiprotons.

Operations reported at 8:27 PM that they had resumed sending beam to NuMI, Pbar, and MCenter.

At 10:06 PM, Operations reported that they were still trying to get MI to send beam to Pbar and NuMI with little success. An operator called in a MI expert.

Operations reported at 10:43 PM that Pbar had resumed stacking, but at a poor stacking rate. Pbar experts were investigating.

Operations reported at 11:11 PM that MI was operating normally. The expert found and fixed a problem module in the Time Line Generator (TLG). Pbar resumed stacking, but again it was at a poor rate. Pbar experts continued their investigations.

Saturday September 17
The midnight shift began with Operations monitoring store 4388 and a small stash, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations reported at 12:32 AM that stacking had returned to normal.

Operations terminated store 4388 at 9:01 AM. Operators prepared the CDF and D0 collision halls for access.

Cryo system technicians began work on wet engines around the Tevatron (TeV) at 9:14 AM.

CDF came out of access around 10:38 AM.

D0 came out of access at 12:20 PM.

Operations put the TeV through a dry squeeze at 12:46 PM.

Operations put the TeV into shot setup at 1:12 PM.

Operations established combination store 4390 at 3:05 PM with an initial luminosity of 102.57E30. Pbar and MCenter resumed normal operations.

At 6:24 PM, a TeV RF station (TRF4) tripped off on an overcurrent indication. This caused a small loss in luminosity and lifetimes.

After having continuing unresolved problems with Booster intensity, Operations called in Booster experts at 6:55 PM.

A Booster expert reported at 11:56 PM that he had resolved the Booster problems.

Sunday September 18
The midnight shift began with Operations monitoring store 4390, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Operations began preparing the Recycler for stashing at 12:11 AM. The stash ended at 1:19 AM and Pbar and MCenter resumed normal operations.

Recycler was prepared for stashing at 5:57 AM. The stash ended at 6:58 AM and Pbar and MCenter resumed normal operations.

At 7:22 AM, Operations reported that Pbar lost its stack due to a ground fault indication (A:IB). The power supply reset, but an operator called in an expert to check the supply.

At 9:05 AM, Operations reported that the Linac intensity had been drifting down. A Linac expert was here and investigating.

At 2:52 PM, the Linac Buncher tripped off. Operators had it back up within nine minutes.

Operations terminated store 4390 at 4:14 PM.

Operations had the D0 collision hall ready for access by 4:30 PM.

At 4:32 PM, Operations reported a TeV sector C3 dump failure indication.

D0 came out of access at 5:19 PM. Operators were still investigating the C3 dump failure. Experts were called in.

At 6:11 PM, Operations reported that a Linac RF station (LRF3) tripped off. Operators investigated several possible problems and then identified the power amplifier filament as the source. An expert was called.

Operators reported that the TeV sector C3 monitor thought there was a dump failure, but there was no indication of an actual failure. While changing out the monitor a pin fell off a cable termination. Operations called in an expert at 7:25 PM to re-terminate the cable.

An expert re-terminated the cable and had the TeV sector C3 chassis volt monitor changed by 7:56 PM. Operations proceeded with the TeV turn on.

The Linac came back on around 8 PM.

Pbar resumed stacking by 8:22 PM. Beam to MCenter, NuMI, and MiniBooNE followed shortly.

At 8:31 PM, LRF3 tripped off for the first of nine times in three and a half hours. It reset each time.

Operations put the TeV through a dry squeeze at 9:16 PM.

Operations put the TeV into shot setup at 9:40 PM.

Operations established combination store 4392 11:31 PM with an initial luminosity of 104.96E30. Pbar and MCenter resumed normal operations.

Monday September 19
The midnight shift began with Operations monitoring store 4390, with Pbar stacking, and with MiniBooNE, NuMI, and MCenter taking beam.

Recycler began preparing to stash at 4:52 AM. The stash ended at 5:38 AM and Pbar and MCenter resumed normal operations.

The Plans for Monday
The plans for today are to stack, stash, and store. The Run coordinator will allow Pbar experts to conduct a reverse proton study later today and TeV experts to conduct a study early Tuesday morning, around 5 AM.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 9/19/2005   email Fermilab